John intends to invest a certain sum of money today in order to earn a return of $6 000 in five years time. how much should john invest today at the interest rate of 7%, compounded annually? would you advise john to proceed with the transction or not? kindly motivate your position.
1
Expert's answer
2013-03-29T06:05:18-0400
FV = PV*(1 + i)^n, where FV - future value, PV - present value, i - interest rate, n - number of years.
In our case, FV = PV + 6000, so we get:
PV + 6000 = PV*1.07^5
1.4026PV - PV = 6000
0.4026PV = 6000
PV = 6000/0.4026 = $14,904.92
We advice John to proceed with this transaction, as he will gain the significant profit. But if he has another more attractive offer, it's better not to proceed with this.